# Bound-state confinement after trap-expansion dynamics in integrable systems

### v1

@@ Abstract
Integrable systems possess stable families of quasiparticles, which are composite objects (bound states) of elementary excitations. Motivated by recent quantum computer experiments, we investigate bound-state transport in the spin-$1/2$ anisotropic Heisenberg chain ($XXZ$ chain). Specifically, we consider the sudden vacuum expansion of a finite region $A$ prepared in a non-equilibrium state. In the hydrodynamic regime, if interactions are strong enough, bound states remain confined in the initial region. Bound-state confinement persists until the density of unbound excitations remains finite in the bulk of $A$. Since region $A$ is finite, at asymptotically long times bound states are “liberated” after the “evaporation” of all the unbound excitations. Fingerprints of confinement are visible in the space-time profiles of local spin-projection operators. To be specific, here we focus on the expansion of the $p$-Néel states, which are obtained by repetition of a unit cell with $p$ up spins followed by $p$ down spins. Upon increasing $p$, the bound-state content is enhanced. In the limit $p\to\infty$ one obtains the domain-wall initial state. We show that for $p<4$, only bound states with $n>p$ are confined at large chain anisotropy. For $p\gtrsim 4$, also bound states with $n=p$ are confined, consistent with the absence of transport in the limit $p\to\infty$. The scenario of bound-state confinement leads to a hierarchy of timescales at which bound states of different sizes are liberated, which is also reflected in the dynamics of the von Neumann entropy.
@@ 1 Introduction
Integrable quantum many-body systems provide an ideal playground to discover exotic behaviors both at equilibrium and out-of-equilibrium. As recognized already by Bethe [1], one of the hallmarks of interacting many-body systems is the formation of multiparticle bound states. In spin chains they are formed by elementary magnon-like excitations. Bound states play a pivotal role in the equilibrium and out-of-equilibrium integrable systems, as they enter prominently in the Thermodynamic Bethe Ansatz [2] approach. Since the birth of Bethe Ansatz, bound states attracted growing attention, also experimentally [3, 4, 5, 6, 7, 8]. Furthermore, bound states are key ingredients behind exotic transport behaviors, such as superdiffusion [9]. Very recently, they have been observed in quantum-computing platforms [10].
Here we investigate the dynamics of bound states after trap-expansion experiments in the Heisenberg spin chain, which is a prototypical integrable system [2]. Our out-of-equilibrium protocol is depicted in Fig. 1: A region $A$ of size $\ell$ is embedded in an infinite chain, and is prepared in a product state. The rest of the chain is in the ferromagnetic state with all the spins down. The ferromagnetic state is the vacuum state for the Bethe Ansatz quasiparticles. The full chain then evolves under the $XXZ$ Hamiltonian. Crucially, the $XXZ$ chain, for anisotropy $\Delta\geq 1$ possesses bound states of an arbitrary number of elementary excitations. During the dynamics, as the state of region $A$ is not an eigenstate of the model, region $A$ undergoes a nontrivial evolution, whereas the remaining part of the chain does not evolve because the ferromagnetic state is eigenstate of the $XXZ$ Hamiltonian. The dynamics of region $A$ is understood in terms of the thermodynamic excitations [2] of the model. In the limit $\ell\to\infty$, at long times local observables in the bulk of region $A$ are described by a thermodynamic macrostate, which is a Generalized Gibbs Ensemble [11] ($GGE$). The initial inhomogeneity that is present at the interface between region $A$ and the rest (see Fig. 1) “melts” during the dynamics. This is due to excitations leaving region $A$, where they are initially produced. At long times and large distances from the interface, the state of the system is described by a space-time dependent $GGE$, which can be obtained within the Generalized Hydrodynamics ($GHD$) approach [12, 13]. In contrast with free-fermion systems, the thermodynamic excitations of the $XXZ$ chain are interacting. However, only elastic scattering is allowed, due to integrability. During the scattering the identity of the quasiparticles is preserved. The net effect of the scatterings is to dress the “bare” velocities of the quasiparticles. The dressed velocities depend on the space-time density of quasiparticles, and hence on the initial state. Crucially, the dressing can be so dramatic that the sign of the velocities change. Typically, this is more striking for the bound states, because their bare velocities (see section 5) are typically smaller than the magnons, implying that the bound states are more affected by scatterings. As we will show, this leads to bound-state confinement, meaning that the bound states remain trapped in region $A$ where they are produced (see Fig. 1). This was observed already in Ref. [14] (see also [15]). Moreover, the sign change of the bound-state velocities was analyzed at the level of the space-time trajectories in Ref. [16], and it has striking effects on the entanglement dynamics [16, 17]. Besides confinement, the presence of bound states gives rise to several intriguing transport behaviors, such as nonanaliticities in the profile of local observables [18].
@@
Here by combining the time-dependent Density Matrix Renormalization Group [19] ($tDMRG$) with state-of-the-art integrability methods [20, 21], we provide a thorough investigation of bound-state confinement in the $XXZ$ chain at anisotropy $\Delta>1$. We show that for generic initial states of region $A$, multiparticle bound states remain confined in $A$. Precisely, this happens provided that interactions, i.e., the anisotropy $\Delta$, are strong enough, and there is a finite density of magnon-like quasiparticles, i.e., unbound ones, in region $A$. Confinement is present at large $\Delta$, and it can be understood perturbatively in the limit $\Delta\to\infty$. Upon lowering $\Delta$ there is a threshold below which the lighter bound states, i.e., formed by two elementary quasiparticles, are not confined. This means that two-particle bound states are allowed to leave region $A$ (see Fig. 1). This is in agreement with the results of Ref. [14] and [15] for the case of Néel and Majumdar-Ghosh initial states. Moreover, for typical initial states of region $A$ there is a hierarchical emission of bound states, meaning that heavier bound states are “liberated” at lower $\Delta$. While confinement can be understood perturbatively, the de-confinement happens near $\Delta=1$, i.e., away from the perturbative regime. This means that the deconfinement pattern of the different bound states depends on the details of the initial state. Finally, it is important to stress that bound-state confinement happens at ballistic level, i.e., in the hydrodynamic limit $x,t\to\infty$ with $\zeta:=x/t$, with $x$ the distance from the interface between $A$ and $B$ (see Fig. 1).
Crucially, since the size $\ell$ of $A$ (see Fig. 1) is finite, at long times tmuch-greater-$t\gg\ell$ the density of magnon quasiparticles in $A$ vanishes. In turn, this implies that after the “evaporation” of all the magnons from region $A$, the system cannot support bound-state confinement. Thus, even for large $\Delta$ bound states cannot remain confined at arbitrarily long times (see Fig. 1 (c)) if region $A$ is finite. Still, the short-time behavior of the spin projectors $P_{\uparrow^{\otimes n}}$ on $n$ consecutive sites are reminiscent of confinement. Following Ref. [22], we use $P_{\uparrow^{\otimes n}}$ as witness for $n$-particles bound-states. We observe that if interactions are not strong enough to support confinement in the hydrodynamic limit, at short times the projectors exhibit a “halo” signal around the interface between $A$ and its complement. On the other hand, if confinement is present, the projectors exhibit a “bottleneck” at the interface between $A$ and $B$ before the “deconfinement” of the bound states happens at longer times.
Finally, we confirm the confinement scenario by considering a “cut” quench (see Fig. 1 (b)) in which the expansion is stopped after a time $t^{*}$. The stopping time is chosen such that the density of magnons in $A$ is finite, i.e., before the bound-state deconfinement starts. For $t>t^{*}$ only part $B$ of the chain is further evolved. If interactions are strong enough to permit bound-state confinement, the projectors $P_{\uparrow^{\otimes n}}$ do not show any signal at $t>t^{*}$ in region $B$.
@@
In this work we focus on the vacuum expansion of the $p$-Néel states, which are “fattened” version of the Néel state. They are defined by repetition of a unit cell containing $p$ up spins followed by $p$ down spins (see Fig. 1 (a)). In the limit $p\to\infty$ the $p$-Néel states converge to the domain-wall state. Interestingly, for $\Delta>1$ transport is suppressed after the quench from the domain-wall state [23]. At $\Delta=1$ transport is neither ballistic nor diffusive, but a superdiffusive dynamics is expected [24]. An interesting feature of the $p$-Néel state is that by changing $p$, one can change the bound-state content. Precisely, states with larger $p$ contain more bound states. For $p=1$ one recovers the standard Néel state, which is somewhat special because for $p=1$ the $GGE$ describing the bulk of $A$ can be determined analytically [25]. This is not the case for $p>2$ [26], although the $GGE$ can be obtained at modest numerical cost by using the results of Ref. [20]. If region $A$ (see Fig. 1) is prepared in the Néel state (with $p=1$), at large enough $\Delta\gtrsim 1.4$, bound states are confined, as observed in Ref. [15]. Interestingly, for $2\leq p\lesssim 4$, only bound states with size $n>p$ are confined, whereas for $p\gtrsim 4$ bound states of size $n=p$ are also confined. Since bound states with $p=n$ dominate the behavior of local observables after the quench from the $p$-Néel states, this is compatible with the absence of transport in the $p\to\infty$ limit.
@@
Finally, we show that bound-state confinement has striking consequences on the dynamics of entanglement entropy. First, within the quasiparticle picture for entanglement dynamics [27, 28, 29], entanglement is created in region $A$. Transport between $A$ and the rest does not generate new entanglement, but entanglement is simply “transported” from $A$ to the rest [17]. We focus on the full expansion of the Néel state. We numerically show that the hierarchical emission of bound states of different sizes is reflected in a multistage growth of the entanglement entropy.
@@
The manuscript is organized as follows. In section 2 we introduce the $XXZ$ spin chain and the out-of-equilibrium protocols. In section 3 we review the Thermodynamic Bethe Ansatz of the model. In section 4 we discuss the Generalized Hydrodynamics ($GHD$) approach. In section 5 by using the $GHD$ approach, we investigate bound-state confinement. In particular, in section 5.1 we provide numerical $GHD$ results confirming bound-state confinement after the expansion of the $p$-Néel states. In section 6, we discuss numerical $tDMRG$ results for the trap-expansion dynamics. In particular, in section 6.1 we discuss the validity of the $GHD$ approach. In section 6.2 and section 6.3 we provide $tDMRG$ data for the full-expansion dynamics and the cut protocol (see Fig. 1 (a) and (b)), respectively. In section 7, we focus on the entanglement entropy after the full-expansion protocol. We conclude and discuss future directions in section 8.
@@ 2 Model and quantum transport protocols
$|N_{p}\rangle:=|\uparrow^{\otimes p}\downarrow^{\otimes p}\rangle^{\otimes L/(2p)},$ (2)
@@
As anticipated, in both protocols $(a)$ and $(b)$ of Fig. 1, at $t>0$ part $A$ of the chain “melts” because $|N_{p}\rangle$ is not an eigenstate of (1). First, the dynamics of an infinite homogeneous chain prepared in $|N_{p}\rangle$ gives rise to a Generalized Gibbs Ensemble [11] ($GGE$) with extensive thermodynamic entropy. Provided that $\ell$ is large enough, local properties in the bulk of region $A$ are described by the $GGE$ arising in the homogeneous chain. Oppositely, at the interface between region $A$ and the rest, the dynamics leads to the formation of a nontrivial profile. Specifically, in the long-time limit, provided that tmuch-less-$t\ll\ell$, to ensure that the central region has a finite density of excitations, ballistic transport sets in. This means that expectation values of generic local observables ${\mathcal{O}}(x,t)$ (with $x$ measured from the interface between $A$ and rest) are functions of $\zeta:=x/t$, and not of space and time separately. The expectation values of ${\mathcal{O}}(\zeta)$ can be obtained (see section 4) by using the Generalized Hydrodynamics [13, 12] ($GHD$). Crucially, since the $XXZ$ chain is interacting, for sufficiently large $\Delta>1$ transport of bound states (of arbitrary size) of elementary excitations from region $A$ to $B$ is inhibited in the ballistic regime. This means that the expectation values ${\mathcal{O}}(\zeta)$ for $\zeta>0$ are fully determined by magnon-like excitations. This bound-state confinement happens for both the protocols of Fig. 1. As already stressed, for the protocol in Fig. 1 $(a)$ the dynamics at asymptotically long times leads to zero density of magnon-like excitations in the bulk of $A$. As we will discuss in section 5 this will lead to the deconfinement of bound states.

### v2

@@ Abstract
Integrable systems possess stable families of quasiparticles, which are composite objects (bound states) of elementary excitations. Motivated by recent quantum computer experiments, we investigate bound-state transport in the spin-$1/2$ anisotropic Heisenberg chain ($XXZ$ chain). Specifically, we consider the sudden vacuum expansion of a finite region $A$ prepared in a non-equilibrium state. In the hydrodynamic regime, if interactions are strong enough, bound states remain confined in the initial region. Bound-state confinement persists until the density of unbound excitations remains finite in the bulk of $A$. Since region $A$ is finite, at asymptotically long times bound states are “liberated” after the “evaporation” of the unbound excitations. Fingerprints of confinement are visible in the space-time profiles of local spin-projection operators. To be specific, here we focus on the expansion of the $p$-Néel states, which are obtained by repetition of a unit cell with $p$ up spins followed by $p$ down spins. Upon increasing $p$, the bound-state content is enhanced. In the limit $p\to\infty$ one obtains the domain-wall initial state. We show that for $p<4$, only bound states with $n>p$ are confined at large chain anisotropy. For $p\gtrsim 4$, also bound states with $n=p$ are confined, consistent with the absence of transport in the limit $p\to\infty$. The scenario of bound-state confinement leads to a hierarchy of timescales at which bound states of different sizes are liberated, which is also reflected in the dynamics of the von Neumann entropy.
@@ 1 Introduction
Integrable quantum many-body systems provide an ideal playground to discover exotic behaviors both at equilibrium and out-of-equilibrium. As recognized already by Bethe [1], one of the hallmarks of interacting many-body systems is the formation of stable multiparticle bound states. Bound states play a pivotal role in the equilibrium and out-of-equilibrium integrable systems, as they enter prominently in the Thermodynamic Bethe Ansatz [2] approach. Since the birth of Bethe Ansatz, bound states attracted growing attention, also experimentally [3, 4, 5, 6, 7, 8]. Furthermore, bound states are key ingredients behind exotic transport behaviors, such as superdiffusion [9]. Very recently, they have been observed in quantum-computing platforms [10, 11]. We should mention that while bound states should be present also in generic, i.e., non integrable models, integrability ensures that they are stable excitations.
Here we investigate the dynamics of bound states after quenches from inhomogeneous initial states in the Heisenberg $XXZ$ spin chain, which is a prototypical integrable system [2]. An important property of integrable systems is that they possess stable quasiparticles, which are labelled by a set of complex numbers called rapidities [2]. For instance, the quasiparticles of the Heisenberg spin chain resemble standard magnon excitations and bound states of magnons. In contrast with standard magnons, however, the excitations of Bethe Ansatz systems are interacting. Crucially, quasiparticles and their rapidities are preserved during the dynamics because integrability allows only elastic scattering. This property allows to understand the dynamics of local observables, and of entanglement [12], in terms of the ballistic propagation of the quasiparticles. This observation is at the heart of the so-called Generalized Hydrodynamics ($GHD$) approach [13, 14].
Our out-of-equilibrium protocol is depicted in Fig. 1: A region $A$ of size $\ell$ is embedded in an infinite chain, and is prepared in a product state. The rest of the chain is in the ferromagnetic state with all the spins down. The ferromagnetic state is the vacuum state for the Bethe Ansatz quasiparticles. The full chain then evolves under the $XXZ$ Hamiltonian. Since quasiparticles are initially present only in region $A$, whereas the remaining part of the chain is in the vacuum state, our protocol is akin to the trap-expansion protocol routinely employed in cold atom experiments. For this reason, in the following we often refer to the protocol in Fig. 1 as trap-expansion or expansion protocol. The initial state is not globally an eigenstate of the $XXZ$ model, implying that the system evolves with time. Specifically, the bulk of region $A$ undergoes a nontrivial evolution from time $t=0^{+}$, whereas the bulk of the remaining part of the chain initially does not evolve, because it is in the vacuum state.
The dynamics of region $A$ is understood in terms of the quasiparticle excitations [2] of the model. For times tmuch-less-$t\ll\ell$, provided that $t$ is long enough for the system to reach local equilibrium, local observables in the bulk of $A$ are described by a thermodynamic macrostate, which is a Generalized Gibbs Ensemble [15] ($GGE$). Moreover, the initial inhomogeneity that is present at the interface between region $A$ and the rest (see Fig. 1) “melts” during the dynamics. This is due to the fact that the quasiparticles that initially are present only in $A$ spread in the rest of the chain. At long times, different regions around the edges of $A$ reach different equilibrium macrostates, giving rise to a space-time dependent $GGE$. This inhomogeneous $GGE$ can be obtained within the $GHD$ approach [13, 14]. The main idea of $GHD$ (see section 4) is that the space-time dependent densities of quasiparticles satisfy a continuity equation, reflecting that quasiparticles do not decay and preserve their rapidity. Moreover, since the scattering between different quasiparticles is elastic, its net effect is to dress the “bare” velocities of the quasiparticles. In turn, the dressed velocities, which appear in the continuity equation, depend on the space-time-dependent densities. This means that the continuity equation has to be solved self-consistently.
Crucially, the effect of the scatterings can be so dramatic that the sign of the velocities changes. Typically, this is more striking for the bound states, because their bare velocities (see section 5) are generically smaller than those of the magnons. As we will show, this leads to bound-state confinement, meaning that the transport of bound states from region $A$ to region $B$ is completely suppressed at ballistic level, i.e. at long times $t$ and large distances $x$ from the interface between $A$ and the rest, with the ratio $\zeta:=x/t$ fixed. In the rest of the paper, the confinement of bound states must always be intended in this limit.
Intuitively, as the bound states approach the interface between $A$ and the rest, they scatter with the magnons, and, as a consequence, they are “pushed back” in $A$. This leads to an accumulation of bound states in $A$. From now on we refer to this behavior, i.e., to the fact that the bound states remain trapped in region $A$ at ballistic level, as bound state confinement. This was observed already in Ref. [16] (see also [17]). Moreover, the sign change of the bound-state velocities was analyzed at the level of the space-time trajectories in Ref. [18], and it has striking effects on the entanglement dynamics [18, 19]. Besides confinement, the presence of bound states gives rise to several intriguing transport behaviors, such as nonanaliticities in the profile of local observables [20].
@@
Finally, we should mention that bound-state confinement gives rise to the so-called quantum distillation [21, 22, 23]. It has been observed in Ref. [21] that after the sudden expansion of interacting fermions if the initial state has a strong admixture of double occupancies, the doublons group together forming a band-insulating state trapped at the center of the system. This state has low-entropy. This protocol has been investigated in cold-atom experiments in [22, 23].
@@
In this work we provide a thorough investigation of bound-state confinement in the $XXZ$ chain at anisotropy $\Delta>1$ by combining the time-dependent Density Matrix Renormalization Group [24] ($tDMRG$) with state-of-the-art integrability methods [25, 26]. For $\Delta>1$ the $XXZ$ chain possesses bound states of an arbitrary number of magnons. We show that all the multiparticle bound states remain confined in $A$, provided that interactions, i.e., the anisotropy $\Delta$, are strong enough, and there is a finite density of magnonic quasiparticles, i.e., unbound ones, in region $A$. Confinement is present at large $\Delta$, and it can be understood perturbatively in the limit $\Delta\to\infty$. Upon lowering $\Delta$ there is a threshold below which the lighter bound states, i.e., formed by two magnons, are not confined. This means that two-particle bound states are allowed to leave region $A$ (see Fig. 1). This is in agreement with the results of Ref. [16] and [17] for the specific cases of Néel and Majumdar-Ghosh initial states. Moreover, for typical initial states of region $A$ there is a hierarchical emission of bound states, meaning that heavier bound states are “liberated” at lower $\Delta$. While confinement can be understood perturbatively, the deconfinement happens near $\Delta=1$, i.e., away from the perturbative regime. This means that the deconfinement pattern of the different bound states depends on the details of the initial state, and it cannot be addressed perturbatively. Moreover, it is important to stress that if the size $\ell$ of region $A$ is finite, confinement can be observed only at intermediate times tmuch-less-$t\ll\ell$. As it will explained in Section 5, intermediate time tmuch-less-$t\ll\ell$ are both a requirement for the applicability of the bipartite protocol in GHD and to ensure that the central region has a finite density of excitations. Indeed, at times tsimilar-$t\sim\ell$ the density of magnons in the bulk of $A$ starts to be depleted. This implies that the effect of the scattering between magnons and bound states, and hence the dressing of bound-state velocity, weakens. At long-enough times, after the magnons “evaporated” from $A$, deconfinement of the bound states happens (see Fig. 1 (c)). Following Ref. [27], we employ the space-time profiles of the spin projectors $P_{x,\uparrow^{\otimes n}}$ to witness bound-state confinement. $P_{x,\uparrow^{\otimes n}}$ are sensitive to the presence of a string of $n$ up spins on consecutive sites starting from $x$. We also confirm the confinement scenario by considering a “cut” quench (see Fig. 1 (b)) in which the expansion is stopped after a time $t^{*}$, and for $t>t^{*}$ only part $B$ of the chain evolves. The stopping time is chosen such that the density of magnons in $A$ is finite, i.e., before the bound-state deconfinement starts. If interactions are strong enough to permit bound-state confinement, the expectation values of the projectors $P_{\uparrow^{\otimes n}}$ do not show any signal at any $t>t^{*}$ in region $B$, since there are no bound states outside of $A$.
@@
In this work, we focus on the expansion of the $p$-Néel states, which are “fattened” version of the Néel state. They are defined by repetition of a unit cell containing $p$ up spins followed by $p$ down spins (see Fig. 1 (a)). An interesting feature of the $p$-Néel state is that by changing $p$, one can change the bound-state content. Precisely, states with larger $p$ contain larger bound states. We observe the following pattern for bound state confinement. If region $A$ (see Fig. 1) is prepared in the Néel state (with $p=1$), at large enough $\Delta\gtrsim 1.4$, all the bound states are confined, as observed in Ref. [17]. Interestingly, for $42\leq p<$ 4, only bound states with size $n>p$ are confined, whereas for $p\geq 4$ bound states of size $n=p$ are also confined.
@@
Finally, we show that bound-state confinement and deconfinement are sharply reflected in the dynamics of the entanglement entropy. First, within the quasiparticle picture for entanglement dynamics [28, 29, 12], entanglement is created in region $A$. Transport between $A$ and the rest does not generate new entanglement, but entanglement is simply “transported” from $A$ to the rest [19]. We focus on the full expansion of the Néel state. We numerically show that the hierarchical emission of bound states of different sizes is reflected in a multistage growth of the entanglement entropy.
@@
The manuscript is organized as follows. In section 2 we introduce the $XXZ$ spin chain and the out-of-equilibrium protocols. In section 3 we review the Thermodynamic Bethe Ansatz ($TBA$) of the model. In section 4 we discuss the Generalized Hydrodynamics ($GHD$) approach. In section 5 by using the $GHD$ approach, we investigate bound-state confinement. In particular, in section 5.1 we provide numerical $GHD$ results confirming bound-state confinement after the expansion of the $p$-Néel states. In section 6, we discuss numerical $tDMRG$ results for the expansion dynamics. In particular, in section 6.1 we discuss the validity of the $GHD$ approach. In section 6.2 and section 6.3 we provide $tDMRG$ data for the full-expansion dynamics and the cut protocol (see Fig. 1 (a) and (b)), respectively. In section 7, we focus on the entanglement entropy after the full-expansion protocol. We conclude and discuss future directions in section 8. In A we discuss the $TBA$ derivation of the $GGE$ for homogeneous quenches. In B we discuss the convergence of the $tDMRG$ method.
@@ 2 Model and quantum transport protocols
$|N_{p}\rangle:=|\uparrow^{\otimes p}\downarrow^{\otimes p}\rangle^{\otimes\ell/(2p)},$ (2)
@@
In the limit $p\to\infty$ the $p$-Néel states converge to the domain-wall state. Interestingly, for $\Delta>1$ transport is suppressed after the quench from the domain-wall state [30]. At $\Delta=1$ transport is neither ballistic nor diffusive, but a superdiffusive dynamics is expected [31].
@@
As anticipated, in both protocols $(a)$ and $(b)$ of Fig. 1, since $|N_{p}\rangle$ is not an eigenstate of (1), the part $A$ of the chain starts to “melt” from arbitrarily small times $t=\epsilon>0$. First, the dynamics of an infinite homogeneous chain prepared in $|N_{p}\rangle$ gives rise to a Generalized Gibbs Ensemble [15] ($GGE$) with extensive thermodynamic entropy. Provided that $\ell$ is large enough, local properties in the bulk of region $A$ are described by the $GGE$ arising in the homogeneous chain. Oppositely, at the interface between region $A$ and the rest, the dynamics leads to the formation of a nontrivial profile. Specifically, in the long-time limit ballistic transport sets in. As it will be explained in Section 5, provided that tmuch-less-$t\ll\ell$ to ensure that the central region has a finite density of excitations, $GHD$ description is valid. This means that expectation values of generic local observables ${\mathcal{O}}(x,t)$ (with $x$ measured from the interface between $A$ and rest) are functions of $\zeta:=x/t$, and not of space and time separately. The expectation values of ${\mathcal{O}}(\zeta)$ can be obtained (see section 4) by using the Generalized Hydrodynamics [14, 13] ($GHD$). Crucially, since the $XXZ$ chain is interacting, for sufficiently large $\Delta>1$ transport of bound states (of arbitrary size) of elementary excitations from region $A$ to $B$ is inhibited in the ballistic regime. This means that the expectation values ${\mathcal{O}}(\zeta)$ for $\zeta>0$ are fully determined by magnonic excitations. This bound-state confinement happens for both the protocols of Fig. 1 at times tmuch-less-$t\ll\ell$. As already stressed, for the protocol in Fig. 1 $(a)$ the dynamics at asymptotically long times leads to zero density of magnonic excitations in the bulk of $A$. As we will discuss in section 5 this will lead to the deconfinement of bound states.